### v2

@@ 4.2 Shortest-path games
We now show the converse implication. Let $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ be a strategy profile such that all players follow $\pi$, and if $\mathcal{P}_{i}$ deviates from $\pi$, the coalition $\mathcal{P}_{-i}$ switches to a winning strategy in the reachability game $(\mathcal{A}_{i},\mathsf{Reach}(T_{i}))$ if $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$ and otherwise the coalition $\mathcal{P}_{-i}$ switches to a strategy that ensures $\min\{\mathsf{val}_{i}(v_{\ell}),\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi{\ell})+1\}$ (we specify a minimum to ensure that the threshold to be ensured is finite) from $v_{\ell}$ if the deviation occurs in $v_{\ell}$. It is immediate that no player has a profitable deviation from $\sigma$ starting in $v_{0}$ thanks to i and ii.
@@ 5 Finite-memory Nash equilibria in reachability and shortest-path games
In this section, we study reachability and shortest-path games (with a single non-negative integer weight function for all players). The main results of this section state that given an NE in a reachability or shortest-path game, we can derive from it a finite-memory NE in which the same targets are visited, the costs are not increased for any player (in shortest-path games) and the memory size of the strategies of the finite-memory NE can be bounded by a function of the number of players. In other words, we obtain arena-independent upper bounds on the sufficient memory for solutions to the constrained NE existence problem (see Section 1). To obtain these arena-independent memory bounds, we do not fully implement the punishment mechanism as in the proofs of Theorems 4.1 and 4.5. Instead, intuitively, following a deviation of $\mathcal{P}_{i}$, the coalition $\mathcal{P}_{-i}$ does not necessarily switch to a punishing strategy for $\mathcal{P}_{i}$. Instead, they may attempt to keep following a suffix of the equilibrium’s original outcome if the deviation does not appear to prevent it.
This section is structured as follows. We illustrate the constructions for reachability and shortest-path games with examples in Section 5.1. In Section 5.2, we show that we can derive, from any NE outcome, another with a simple structure. We build on such NE outcomes to define finite-memory NE outcomes. In Section 5.3, we describe the base of the Mealy machine structure that we use to construct finite-memory NEs in reachability and shortest-path games. Sections 5.4 and 5.5 explain how to extend the Mealy machines of Section 5.3 to construct finite-memory NE in reachability and shortest-path games respectively.
@@ 5.1 Examples
In this section, we illustrate the upcoming construction for finite-memory NEs for both settings of interest. We start with a reachability game.
@@
We consider the game on the arena depicted in Figure 3(a) where the objective of $\mathcal{P}_{i}$ is $\mathsf{Reach}(\{t_{i}\})$ for all $i\in\llbracket{}4\rrbracket{}$. This game has a similar structure to that of the game of Example 2.3. To illustrate the idea behind the upcoming construction, we present a finite-memory NE from $v_{0}$ in which the targets of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are visited in its outcome. We define NEs from well-structured NE outcomes: we consider the play $\pi=v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$ in this example. Intuitively, this is the simplest NE outcome starting in $v_{0}$ that visits both $t_{1}$ and $t_{2}$.
[…]
(a) An arena. Circles, squares, diamonds and hexagons are resp. $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$ vertices.
@@
First, observe that $\pi$ can be seen as the combination of the simple history $\mathsf{sg}_{1}=v_{0}v_{1}v_{2}t_{1}$ and the simple lasso $\mathsf{sg}_{2}=t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$. The simple history $\mathsf{sg}_{1}$ connects the initial vertex to the first visited target, and the simple lasso $\mathsf{sg}_{2}$ connects the first target to the second and contains the suffix of the play. Therefore, if we were not concerned with the stability of the equilibrium, the outcome $\pi$ could be obtained by using a finite-memory strategy profile where all strategies are defined by a Mealy machine with state space $\llbracket{}2\rrbracket{}$. Intuitively, these strategies would follow $\mathsf{sg}_{1}$ while remaining in their first memory state $1$, then, when $t_{1}$ is visited, they would update their memory state to $2$ and follow $\mathsf{sg}_{2}$.
@@
By construction, if $\mathcal{P}_{i}$ (among $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$) deviates and exits the set of vertices of $\mathsf{sg}_{j}$ when in a memory state of the form $(\cdot,j)$, then the memory updates to $(\mathcal{P}_{i},j)$ and does not change until the play returns to some vertex of $\mathsf{sg}_{j}$ (which is not possible here due to the structure of the arena, but may be in general). For instance, assume $\mathcal{P}_{3}$ moves from $v_{1}$ to $v_{3}$ after the history $h=v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}$. Then the memory after $h$ is in state $(\mathcal{P}_{3},2)$ and no longer changes from there on.
@@
We consider the shortest-path game on the weighted arena depicted in Figure 4(a) where the target of $\mathcal{P}_{i}$ is $T_{i}=\{t,t_{12}\}$ for all $i\in\llbracket{}2\rrbracket{}$ and $T_{3}=\{t\}$ for $\mathcal{P}_{3}$. Our goal is to construct an NE from $v_{0}$ with the cost profile $(5,5,5)$; this can be done by designing an NE with outcome $\pi=v_{0}v_{1}v_{3}t^{\omega}$. We establish that a finite-memory NE with outcome $\pi$ from $v_{0}$ cannot be obtained with the construction of Example 5.1. We then provide an alternative construction that builds on the same ideas.
@@
(a) A weighted arena. Edge labels indicate their weight. Unlabelled edges have a weight of $1$.
@@ 5.2 Well-structured Nash equilibrium outcomes
The examples of finite-memory NEs presented on the previous section were built from well-structured NE outcomes. In this section, we show that such well-structured NE outcomes can always be found in reachability and shortest-path games. More precisely, we show that we can always derive, from a given NE outcome, another NE outcome that admits a finite simple decomposition such that each history segment of this decomposition connects the different targets visited along the play and has minimal weight. In particular, the derived NE outcome has a preferable cost profile to that of the original outcome. This derivation can be seen as a simplification process, illustrated in Figure 5: we decompose a given NE outcome into segments up connecting the first visit to target sets, and then simplify each segment individually while minimising the weight of the simple history segments.
@@
Figure 5: Simplification process for an NE outcome in a multi-player shortest-path game. Doubly circles vertices denote the first occurrence of a target state of some player in the play.
To formalise the existence of well-structured NE outcomes, we focus on shortest-path games. The result can be recovered for reachability games by noting that a shortest-path game with a constant zero weight function is equivalent to its unweighted reachability game. Our simplification process yields outcomes admitting a decomposition satisfying the properties outlined in the following lemma.
@@
    if $k>0$, then for all $j\in\llbracket{}k\rrbracket{}$, $w(\mathsf{sg}_{j})$ is minimum among all histories that share their first and last vertex with $\mathsf{sg}_{j}$ and traverse a subset of the vertices occurring in $\mathsf{sg}_{j}$; and
@@
We distinguish two cases in this proof. First, let us assume that $k=0$, i.e., $\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\subseteq\{0\}$. If $\pi^{\prime}$ is a simple play, it suffices to choose $\pi=\pi^{\prime}$; it admits the trivial simple decomposition $(\pi)$. Therefore, we assume that $\pi^{\prime}$ is not a simple play. This implies that there is a simple lasso $\pi\in\mathsf{Plays}(\mathcal{A})$ starting from $v_{0}$ that only uses vertices that occur in $\pi^{\prime}$. It follows that $\mathsf{VisPos}_{\mathcal{G}}(\pi)\subseteq\{0\}$. By Theorem 4.5, $\pi$ is an NE outcome; the first condition of the characterisation of Theorem 4.5 follows from it holding for $\pi^{\prime}$ and the second condition holds because $\mathsf{VisPos}_{\mathcal{G}}(\pi)\subseteq\{0\}$.
Second, we assume that $k>0$. We define $\pi$ by describing the simple decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$. Let $\ell_{1}^{\prime}<\ldots<\ell_{k}^{\prime}$ be the elements of $\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\setminus\{0\}$ and $\ell_{0}^{\prime}=0$. For $j\in\llbracket{}k\rrbracket{}$, we let $\mathsf{sg}^{\prime}_{j}$ denote the segment of $\pi$ between positions $\ell_{j-1}^{\prime}$ and $\ell_{j}^{\prime}$. We let $\mathsf{sg}_{j}$ be a simple history that shares its first and last vertex with $\mathsf{sg}_{j}^{\prime}$ and traverses a subset of the vertices occurring in $\mathsf{sg}_{j}^{\prime}$, with minimal weight among all such histories. It remains to define the segment $\mathsf{sg}_{k+1}$. We let $\mathsf{sg}_{k+1}$ be $\pi^{\prime}{\ell_{k}}$ if $\mathsf{sg}_{k}\cdot\pi^{\prime}{\ell_{k}}$ is a simple play, and otherwise we let $\mathsf{sg}_{k+1}$ be any play starting in $\mathsf{last}(\mathsf{sg}_{k})$ such that $\mathsf{sg}_{k}\cdot\mathsf{sg}_{k+1}$ is a simple lasso in which only vertices of $\pi^{\prime}$ occur. It follows from this choice of $\mathsf{sg}_{k+1}$ that $\pi=\mathsf{sg}_{1}\cdot\ldots\cdot\mathsf{sg}_{k+1}$ and $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ satisfy i.
We now prove that the play $\pi$ is an NE outcome satisfying conditions ii-iv. Let $\pi=v_{0}v_{1}\ldots$, $\ell_{1}<\ldots<\ell_{k}$ be the elements of $\mathsf{VisPos}_{\mathcal{G}}(\pi)\setminus\{0\}$ and $\ell_{0}=0$. To show that $\pi$ is an NE outcome, we rely on Theorem 4.5. Because $\pi^{\prime}$ is an NE outcome and all vertices occurring in $\pi$ occur in $\pi^{\prime}$, it follows that the first condition of the characterisation of Theorem 4.5 holds for $\pi$.
@@
We prove that the first inequality holds by contradiction. Assume that $\mathsf{SPath}^{T_{i}}_{w}(\pi{\ell})>\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime}{\ell^{\prime}})$. It must be the case that either a suffix of $\mathsf{sg}^{\prime}_{j}$ starting in $v_{\ell}$ must have weight strictly less than the suffix $v_{\ell}\ldots v_{\ell_{j}}$ of $\mathsf{sg}_{j}$, or that $w(\mathsf{sg}_{j^{\prime}})>w(\mathsf{sg}_{j^{\prime}}^{\prime})$ for some $j<j^{\prime}\leq j_{i}$. Both possibilities contradict the choice of the elements of $\mathcal{D}$, therefore we have $\mathsf{SPath}^{T_{i}}_{w}(\pi{\ell})\leq\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime}{\ell^{\prime}})$. The second inequality holds by Theorem 4.5 as $\pi^{\prime}$ is an NE outcome. We remark (for condition iv) that in the special case $\ell=0$, the first inequality implies that $\mathsf{SPath}^{T_{i}}_{w}(\pi)\leq\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime})$ as we can choose $\ell^{\prime}=0$. We have shown that $\pi$ is an NE outcome.
We now show that conditions ii-iv hold. Condition ii follows immediately by construction. Let $j\in\llbracket{}k\rrbracket{}$. The minimum in condition iii for $j$ is attained by some simple history. By construction, it must be realised by $\mathsf{sg}_{j}$, implying that condition iii holds. For condition iv, due to the above, we need only consider players who do not see their target. For these players, the condition follows from the equality $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$ implying that players have an infinite cost in $\pi$ if and only if they have an infinite cost in $\pi^{\prime}$. This concludes the proof that conditions ii-iv are satisfied by $\pi$.
Condition i of the statement of Lemma 5.3 suggests that we could merge the last two segments of the decomposition provided by the lemma to obtain simple decompositions with one less element. In our upcoming NE construction (see the examples of Section 5.1), players do not react to deviations that remain within the current segment of the intended outcome. In the following example, we illustrate that this behaviour can be exploited by players in shortest-path games if we merge these last two segments. This motivates the separation of these two segments.
@@
Condition i on histories in Lemma 5.3 is relevant for reachability games. In a reachability game, players whose targets are visited cannot have profitable deviations. This implies that the issue in shortest-path games highlighted by Example 5.5 cannot occur in reachability games, i.e., we can merge the last two segments of the decomposition provided by Lemma 5.3 and construct NE from the smaller decomposition. A smaller decomposition allows us to have smaller memory bounds for finite-memory NEs in reachability games.

### v3

@@ 4.2 Shortest-path games
We now show the converse implication. Let $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$ be a strategy profile such that all players follow $\pi$, and if $\mathcal{P}_{i}$ deviates from $\pi$, the coalition $\mathcal{P}_{-i}$ switches to a memoryless strategy $\tau_{-i}$ that is uniformly winning for the second player of the coalition reachability game $(\mathcal{A}_{i},\mathsf{Reach}(T_{i}))$ and that ensures at cost of at most $\min\{\mathsf{val}_{i}(v),\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi)\}$ from all $v\in V$ (such a strategy exists by Theorem 3.7 in $\mathcal{G}_{i}$).
We now show that for all $i\in\llbracket{}n\rrbracket{}$, $\mathcal{P}_{i}$ does not have a profitable deviation. This is equivalent to showing that for all plays $\pi^{\prime}$ starting in $v_{0}$, if $\pi^{\prime}$ is consistent with $\sigma_{-i}$, then $\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime})\geq\mathsf{SPath}^{T_{i}}_{w}(\pi)$. We let $\pi^{\prime}=v_{0}^{\prime}v_{1}^{\prime}\ldots\in\mathsf{Plays}(\mathcal{A})$ such that $v_{0}^{\prime}=v_{0}$ and $\pi^{\prime}$ is consistent with $\sigma_{-i}$. We assume that $\pi^{\prime}\neq\pi$, as the inequality is direct otherwise, and let $r=\min\{\ell\in\mathbb{N}\mid v_{\ell}\neq\ell\}$. By construction of $\sigma$ and choice of $r$, it follows that $\pi^{\prime}{r-1}$ is consistent with the punishing strategy $\tau_{-i}$.
First, we assume that $i\notin\mathsf{SatPl}_{\mathcal{G}}(\pi)$. In this case, we must show that $\pi^{\prime}\notin\mathsf{Reach}(T_{i})$. On the one hand, no elements of $T_{i}$ occur in $\pi^{\prime}{r-1}$ by choice of $\tau_{-i}$, as i yields $v^{\prime}_{r-1}=v_{r-1}\notin W_{i}(\mathsf{Reach}(T_{i}))$. On the other hand, no target vertices of $\mathcal{P}_{i}$ occur in $\pi^{\prime}{r-1}=\pi{r-1}$ as $\pi\notin\mathsf{Reach}(T_{i})$. This ends the proof of the first case.
We now assume that $i\in\mathsf{SatPl}_{\mathcal{G}}(\pi)$. If a target vertex of $\mathcal{P}_{i}$ occurs in $\pi{r-1}$, then $\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi)=\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi^{\prime})$ (which is sufficient to end the proof). We therefore assume that no vertex in $\pi{r-1}$ is in $T_{i}$ and further distinguish two cases. We first assume that $\mathsf{val}_{i}(v_{r-1})\geq\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi)$. We have
$\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi^{\prime})=w(\pi{r-1})+\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi^{\prime}{r-1})\geq\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi),$
as $\tau_{-i}$ ensures a cost of at least $\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi)$ from $v_{r-1}$. We now assume that $\mathsf{val}_{i}(v_{r-1})<\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi)$. In this case, we obtain that
$\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi^{\prime})$ $=w(\pi{r-1})+\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi^{\prime}{r-1})$
$\geq w(\pi{r-1})+\mathsf{val}_{i}(v_{r-1})$
$\geq w(\pi{r-1})+\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi{r-1})$
$=\mathsf{SPath}^{T_{i}}_{w_{i}}(\pi),$
where the penultimate inequality follows from ii.
@@ 5 Finite-memory Nash equilibria in reachability and shortest-path games
In this section, we study reachability and shortest-path games (with a single non-negative integer weight function for all players). The main results of this section state that given an NE in a reachability or shortest-path game, we can derive from it a finite-memory NE in which the same targets are visited, the costs are not increased for any player (in shortest-path games) and we provide an upper bound depending solely on the number of players on the amount of memory used by all of the involved strategies. In other words, we obtain arena-independent upper bounds on the sufficient memory for solutions to the constrained NE existence problem (see Section 1). To obtain these arena-independent memory bounds, we do not fully implement the punishment mechanism as in the proofs of Theorems 4.1 and 4.5. Instead, intuitively, following a deviation of $\mathcal{P}_{i}$, the coalition $\mathcal{P}_{-i}$ does not necessarily switch to a punishing strategy for $\mathcal{P}_{i}$. Instead, they may attempt to keep following a suffix of the equilibrium’s original outcome if the deviation does not appear to prevent it.
This section is structured as follows. We illustrate the constructions for reachability and shortest-path games with examples in Section 5.1. In Section 5.2, we show that we can derive, from any NE outcome, an NE outcome that admits a finite simple decomposition. We impose technical properties on these decompositions such that we can construct finite-memory NEs based on them. We introduce (finite-memory) strategies based on simple play decompositions in Section 5.3. Sections 5.4 and 5.5 explain how to instantiate the Mealy machines of Section 5.3 to construct finite-memory NEs in reachability and shortest-path games respectively.
@@ 5.1 Examples
In this section, we illustrate the upcoming construction for finite-memory NEs for both settings of interest. We construct NEs from (well-chosen) simple segment decompositions of plays. The main role of memory in the following is to keep track of two pieces of information: a segment of the play in the considered decomposition (which describes the moves to be made), and the last player to have moved (to track who to punish if necessary). We start with a reachability game.
@@
We consider the game on the arena depicted in Figure 3(a) where the objective of $\mathcal{P}_{i}$ is $\mathsf{Reach}(\{t_{i}\})$ for all $i\in\llbracket{}4\rrbracket{}$. This game has a similar structure to that of the game of Example 2.3. To illustrate the idea behind the upcoming construction, we present a finite-memory NE from $v_{0}$ in which the targets of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are visited in its outcome. In this example, we construct an NE from the NE outcome $\pi=v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$. Intuitively, this is the simplest NE outcome starting in $v_{0}$ that visits both $t_{1}$ and $t_{2}$.
$v_{0}$$t_{2}$$v_{1}$$v_{2}$$t_{1}$$v_{4}$$v_{3}$$v_{5}$$t_{3}$$t_{4}$
(a) An arena. Circles, squares, diamonds and hexagons are respectively $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, $\mathcal{P}_{3}$, $\mathcal{P}_{4}$ vertices.
@@
First, observe that $\pi$ can be seen as the combination of the simple history $\mathsf{sg}_{1}=v_{0}v_{1}v_{2}t_{1}$ and the simple lasso $\mathsf{sg}_{2}=t_{1}v_{2}v_{1}v_{0}t_{2}^{\omega}$. The simple history $\mathsf{sg}_{1}$ connects the initial vertex to the first visited target, and the simple lasso $\mathsf{sg}_{2}$ connects the first target to the second and contains the suffix of the play. Therefore, if $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$ were not looking for a profitable deviation, the outcome $\pi$ could be obtained by using a finite-memory strategy profile where all strategies are defined by a Mealy machine with state space $\llbracket{}2\rrbracket{}$. Intuitively, these strategies would follow $\mathsf{sg}_{1}$ while remaining in their first memory state $1$, then, when $t_{1}$ is visited, they would update their memory state to $2$ and follow $\mathsf{sg}_{2}$.
@@
By construction, if $\mathcal{P}_{i}$ (among $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$) deviates and exits the set of vertices of $\mathsf{sg}_{j}$ when in a memory state of the form $(\cdot,j)$, then the memory updates to $(\mathcal{P}_{i},j)$ and does not change until the play returns to some vertex of $\mathsf{sg}_{j}$ (which is not possible here due to the structure of the arena, but may be in general). For instance, assume that $\mathcal{P}_{3}$ moves from $v_{1}$ to $v_{4}$ after the history $h=v_{0}v_{1}v_{2}t_{1}v_{2}v_{1}$. Then the memory after $h$ is in state $(\mathcal{P}_{3},2)$ and no longer changes from there on.
@@
We consider the shortest-path game on the weighted arena depicted in Figure 4(a) where the target of $\mathcal{P}_{i}$ is $T_{i}=\{t,t_{12}\}$ for all $i\in\llbracket{}2\rrbracket{}$ and $T_{3}=\{t\}$ for $\mathcal{P}_{3}$. Our goal is to construct an NE from $v_{0}$ with the cost profile $(5,5,5)$; this can be done by designing an NE with outcome $\pi=v_{0}v_{1}v_{3}t^{\omega}$. We argue that the idea described in Example 5.1 would not work in this setting. We then provide an alternative construction that builds on the same ideas.
@@
(a) A weighted arena. Circles, squares and diamonds are respectively $\mathcal{P}_{1}$, $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ vertices. Edge labels indicate their weight. Unlabelled edges have a weight of $1$.
@@ 5.2 Well-structured Nash equilibrium outcomes
The examples of finite-memory NEs presented on the previous section were built from well-chosen decompositions of an NE outcome. In this section, we show that NE outcomes that admit a decomposition from which we can construct an NE can always be found in reachability and shortest-path games. More precisely, we show that we can always derive, from a given NE outcome, another NE outcome that admits a finite simple decomposition such that each history segment of this decomposition connects the different targets visited along the play and has minimal weight. In particular, the derived NE outcome has a preferable cost profile to that of the original outcome.
@@
Figure 5: Simplification process for an NE outcome in a multi-player shortest-path game. Doubly circled vertices denote the first occurrence of a target vertex of some player in the play.
We focus on shortest-path games in this section. For reachability games, the existence of suitable NE outcomes can be recovered by noting that a shortest-path game with a constant zero weight function is equivalent to its unweighted reachability game. This derivation can be seen as a simplification process, illustrated in Figure 5: we decompose a given NE outcome into segments up connecting the first visit to target sets, and then simplify each segment individually while minimising the weight of the simple history segments. This simplification process yields outcomes admitting a decomposition satisfying the properties outlined in the following lemma.
@@
    if $k>0$, then for all $j\in\llbracket{}k\rrbracket{}$, $w(\mathsf{sg}_{j})$ is minimal among all histories that share their first and last vertex with $\mathsf{sg}_{j}$ and traverse a subset of the vertices occurring in $\mathsf{sg}_{j}$; and
@@
We distinguish two cases in this proof. First, let us assume that $k=0$, i.e., $\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\subseteq\{0\}$. If $\pi^{\prime}$ is a simple play, it suffices to choose $\pi=\pi^{\prime}$; it admits the trivial simple decomposition $(\pi)$. Therefore, we assume that $\pi^{\prime}$ is not a simple play. This implies that there is a simple lasso $\pi\in\mathsf{Plays}(\mathcal{A})$ starting from $v_{0}$ that only uses vertices that occur in $\pi^{\prime}$. It follows that $\mathsf{VisPos}_{\mathcal{G}}(\pi)\subseteq\{0\}$. By Theorem 4.5, $\pi$ is an NE outcome; the first condition of the characterisation of Theorem 4.5 follows from it holding for $\pi^{\prime}$ and the second condition holds because $\mathsf{VisPos}_{\mathcal{G}}(\pi)=\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\subseteq\{0\}$.
Second, we assume that $k>0$. We define $\pi$ by describing the simple decomposition $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$. Let $\ell_{1}^{\prime}<\ldots<\ell_{k}^{\prime}$ be the elements of $\mathsf{VisPos}_{\mathcal{G}}(\pi^{\prime})\setminus\{0\}$ and $\ell_{0}^{\prime}=0$. For $j\in\llbracket{}k\rrbracket{}$, we let $\mathsf{sg}^{\prime}_{j}$ denote the segment of $\pi$ between positions $\ell_{j-1}^{\prime}$ and $\ell_{j}^{\prime}$. We let $\mathsf{sg}_{j}$ be a simple history that shares its first and last vertex with $\mathsf{sg}_{j}^{\prime}$ and traverses a subset of the vertices occurring in $\mathsf{sg}_{j}^{\prime}$, with minimal weight among all such histories. This definition ensures that, for all $j\in\llbracket{}k\rrbracket{}$, the minimum in condition iii is realised by $\mathsf{sg}_{j}$, and thus condition iii holds. It remains to define the segment $\mathsf{sg}_{k+1}$. We let $\mathsf{sg}_{k+1}$ be $\pi^{\prime}{\ell_{k}}$ if $\mathsf{sg}_{k}\cdot\pi^{\prime}{\ell_{k}}$ is a simple play, and otherwise we let $\mathsf{sg}_{k+1}$ be any play starting in $\mathsf{last}(\mathsf{sg}_{k})$ such that $\mathsf{sg}_{k}\cdot\mathsf{sg}_{k+1}$ is a simple lasso in which only vertices of $\pi^{\prime}$ occur. It follows from this choice of $\mathsf{sg}_{k+1}$ that $\pi=\mathsf{sg}_{1}\cdot\ldots\cdot\mathsf{sg}_{k+1}$ and $\mathcal{D}=(\mathsf{sg}_{1},\ldots,\mathsf{sg}_{k+1})$ satisfy i.
We now prove that $\pi$ is an NE outcome satisfying ii and iv. Let $\pi=v_{0}v_{1}\ldots$, and let $\ell_{1}<\ldots<\ell_{k}$ be the elements of $\mathsf{VisPos}_{\mathcal{G}}(\pi)\setminus\{0\}$ and $\ell_{0}=0$. To show that $\pi$ is an NE outcome, we rely on Theorem 4.5. Because $\pi^{\prime}$ is an NE outcome and all vertices occurring in $\pi$ occur in $\pi^{\prime}$, it follows that the first condition of the characterisation of Theorem 4.5 holds for $\pi$.
@@
We prove that the first inequality holds by contradiction. Assume that $\mathsf{SPath}^{T_{i}}_{w}(\pi{\ell})>\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime}{\ell^{\prime}})$. It must be the case that either a suffix of $\mathsf{sg}^{\prime}_{j}$ starting in $v_{\ell}$ has weight strictly less than the suffix $v_{\ell}\ldots v_{\ell_{j}}$ of $\mathsf{sg}_{j}$, or that $w(\mathsf{sg}_{j^{\prime}})>w(\mathsf{sg}_{j^{\prime}}^{\prime})$ for some $j<j^{\prime}\leq j_{i}$. Both possibilities contradict the choice of the elements of $\mathcal{D}$, therefore we have $\mathsf{SPath}^{T_{i}}_{w}(\pi{\ell})\leq\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime}{\ell^{\prime}})$. The second inequality holds by Theorem 4.5 as $\pi^{\prime}$ is an NE outcome. We remark (for condition iv) that in the special case $\ell=0$, the first inequality implies that $\mathsf{SPath}^{T_{i}}_{w}(\pi)\leq\mathsf{SPath}^{T_{i}}_{w}(\pi^{\prime})$ as we can choose $\ell^{\prime}=0$. We have shown that $\pi$ is an NE outcome.
We conclude by proving that ii and iv hold. Condition ii follows immediately by construction. Let $j\in\llbracket{}k\rrbracket{}$. For condition iv, due to the above, we need only consider players who do not see their target. For these players, the condition follows from the equality $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$ implying that players have an infinite cost in $\pi$ if and only if they have an infinite cost in $\pi^{\prime}$. This concludes the proof.
Condition i of the statement of Lemma 5.3 suggests that we could merge the last two segments of the decomposition provided by the lemma to obtain simple decompositions with one less element. In reachability games, we can build NEs from these smaller decompositions, and thus obtain smaller Mealy machines than by using the larger decomposition. However, this is not true in shortest-path games. In our upcoming NE construction (see the examples of Section 5.1), players do not react to deviations that remain within the current segment of the intended outcome. In the following example, we illustrate that this behaviour can be exploited by players in shortest-path games if we merge these last two segments.